# OSMa-Bench: Evaluating Open Semantic Mapping Under Varying Lighting Conditions

### v1

@@ V-B Experimental Results
The evaluation was conducted for three methods: ConceptGraphs [gu2024conceptgraphs], OpenScene [peng2023openscene], and a very recent BBQ [linok2024beyond]. These methods employ different approaches, so we can perform benchmarking on a wider spectrum.
Obtained results for segmentation quality metrics for all considered methods and scenes are presented in Tables I-III, while its relative change under different test conditions are illustrated on a ReplicaCAD [szot2021habitat] dataset by Fig. 5.
Semantic graph quality assessment have been performed according to the described methodology. Fig 6 illustrates VQA accuracy achieved by BBQ [linok2024beyond] and ConceptGraphs [gu2024conceptgraphs] methods for different categories of semantic queries.
For BBQ [linok2024beyond], question related to object relations were not considered, as its scene graph construction is query-dependent. Instead, the evaluation focused on assessing the correctness of object presence and attributes, ensuring that the retrieved object set aligns with the given question.
TABLE I: mAcc Values on ReplicaCAD($\uparrow$)
Approach FPS (Hz) Baseline Camera Light Dynamic Lights Nominal Lights Velocity Min Max Avg
ConceptGraphs[gu2024conceptgraphs] $0.19^{*}$ 0.291 0.276 0.260 0.258 0.283 0.258 0.291 0.274
BBQ[linok2024beyond] $0.93^{*}$ 0.162 0.165 0.128 0.157 0.140 0.128 0.165 0.150
Openscene[peng2023openscene] 0.13 0.130 0.127 0.140 0.136 0.132 0.127 0.140 0.133
@@
TABLE II: f-mIoU Values on ReplicaCAD ($\uparrow$)
Approach Baseline Camera Light Dynamic Lights Nominal Lights Velocity Min Max Avg
ConceptGraphs[gu2024conceptgraphs] 0.324 0.296 0.299 0.310 0.309 0.296 0.324 0.308
BBQ[linok2024beyond] 0.238 0.232 0.179 0.205 0.213 0.179 0.238 0.213
Openscene[peng2023openscene] 0.484 0.482 0.479 0.469 0.475 0.469 0.484 0.478
TABLE III: Semantic Segmentation on HM3D
Approach FPS (Hz) Camera Light Nominal Lights Velocity
mAcc $\uparrow$ f-mIoU $\uparrow$ mAcc $\uparrow$ f-mIoU $\uparrow$ mAcc $\uparrow$ f-mIoU $\uparrow$
ConceptGraphs[gu2024conceptgraphs] $0.16^{*}$ 0.143 0.150 0.158 0.169 0.156 0.156
BBQ[linok2024beyond] $0.92^{*}$ 0.085 0.202 0.124 0.207 0.087 0.187
Openscene[peng2023openscene] 0.25 0.116 0.293 0.118 0.352 0.121 0.344
@@
Figure 5: Comparison of degradation for mAcc and f-mIoU across different environmental conditions (less means less robust). The degradation is computed as $D(\%)=\frac{C-B}{B}\times 100$, where $B$ is the baseline metric and $C$ is the metric under the given condition.
Figure 6: Comparison of Answering Accuracy. We evaluate BBQ  [linok2024beyond] and ConceptGraphs [gu2024conceptgraphs] across ReplicaCAD [szot2021habitat] and HM3D [ramakrishnan2021habitat] sequences for different configurations.
In addition, we measure time consumption using NVIDIA GeForce RTX 4090 for 3D scene representation in frames per second (FPS), which is shown in Table  I and Table  III.
@@ V-C Results Analysis
Through this systematic variation and evaluation, we aimed to comprehensively understand the robustness and reliability of SOTA methods under different environmental conditions and configurations. The major results are:
  • ConceptGraphs [gu2024conceptgraphs] showed the best value for the mAcc metric on all datasets, but still struggles to accurately detect small objects.
  • BBQ [linok2024beyond], taking into account the resulted FPS, seemed to be faster than ConceptGraphs [gu2024conceptgraphs], while BBQ [linok2024beyond] has a significant decrease in metric values under different lighting conditions and turned out to be the most sensitive.
  • OpenScene [peng2023openscene] turned out to be the most stable method among those tested (best f-mIoU overall). The advantage of this approach is a rich semantic description of the scene. However, the quality of the constructed semantic masks is quite poor. We assume that it happens due to the usage of pixel-aligned embeddings, not a segmentation model.
Figure 6 presents the answering accuracy across different question categories and scene modifications. Binary existence-based questions exhibit high accuracy, indicating that models do not hallucinate non-existent objects. Overall, binary questions achieve accuracy above 50–60%, confirming non-random response patterns.
@@
The highest accuracy is observed under nominal conditions — the baseline for ReplicaCAD [szot2021habitat] and nominal lighting for HM3D [ramakrishnan2021habitat]. Variations in lighting and velocity reduce accuracy, especially in object attributes and relationships questions. Attribute-based questions often fail due to missing material and property details in scene graph captions, leading to incomplete LLM responses.
Taking into account the analysis of metrics, it was also possible to detect the following corner cases under variable lighting conditions, which can lead to a decrease in accuracy. Examples of the following corner cases are shown in Fig.  7. When using camera light, shadows of objects can be counted as part of the object. Dynamic lighting causes incorrect segmentation of uniform surfaces, such as floors, due to illumination differences. Also, when using only the camera light source, a noisy point cloud was obtained. The tested methods show sensitivity to objects with a standard description (for example, “wall cabinet” and “base cabinet”). Also, when assessing the quality of semantic segmentation, it turned out that the methods show worse results if we use a ‘_’ instead of a space in the class name.
Object merging issues occur, for example in ConceptGraphs on ReplicaCAD, where staircases are split into individual steps. This impacts Measurement-based question answering, e.g.:
@@
Similar errors appear in BBQ, inflating instance counts.
@@ VI Conclusion
In this paper, we target a systematic evaluation of open semantic mapping algorithms, in particular their robustness in dynamic environments, e.g. under varying indoor lighting conditions. To accomplish the task we introduced a dynamically configurable and highly automated LLM/LVLM-powered benchmarking methodology.
Our novel dataset enabled a comprehensive analysis of leading methods, such as OpenScene [peng2023openscene], BBQ [linok2024beyond], and Concept Graph [gu2024conceptgraphs], highlighting the impact of lighting variations on semantic segmentation accuracy and quality of constructed knowledge graphs.
@@
In our future work, we aim to expand the proposed methodology to assess a broader range of OSM methods focusing on skills crucial for autonomous navigation and general-purpose robots in dynamic environments. In particular, benchmarking visual grounding and scene graph evaluation skills will require adapting the suggested VQA pipeline either to generate question based on ground truth info on all instances and its positions (if available from the dataset) or to process a complete scene representation rather than analyzing it frame-by-frame. Verifying actionable affordances, spatial reasoning, and traversability skills will require only extending questions’ categories within the VQA pipeline, i.e. adapting the service prompts. We also did a preliminary study on incorporating synthetic dynamic objects into scenes and will continue to develop this aspect for the upcoming edition of OSMa-Bench.
[…]

### v2

@@ V-B Experimental Results
TABLE I: Performance on ReplicaCAD
Approach FPS (Hz) Baseline Camera Light Dynamic Lights Nominal Lights Velocity
mAcc f-mIoU mAcc f-mIoU mAcc f-mIoU mAcc f-mIoU mAcc f-mIoU
ConceptGraphs [1] $0.19^{*}$ 29.53 28.00 29.76 22.71 24.73 14.29 26.80 21.34 28.32 27.33
BBQ [2] $0.93^{*}$ 25.25 36.42 26.81 29.59 24.60 32.25 25.07 33.33 26.63 32.30
OpenScene [3] 0.13 24.17 51.17 22.82 48.83 22.02 49.05 22.16 47.09 25.01 55.81
@@
TABLE II: Performance on HM3D
Approach FPS (Hz) Nominal Lights Camera Light Velocity
mAcc f-mIoU mAcc f-mIoU mAcc f-mIoU
ConceptGraphs [1] $0.16^{*}$ 20.39 17.72 16.73 15.48 19.42 17.49
BBQ [2] $0.92^{*}$ 18.14 23.96 17.24 24.03 18.28 24.71
OpenScene [3] 0.25 20.58 34.19 17.29 29.97 19.77 34.02
@@
Figure 5: Comparison of changes (%) in mAcc and f-mIoU across conditions relative to the baseline on the ReplicaCAD dataset.
Figure 6: Comparison of Answering Accuracy across ReplicaCAD and HM3D sequences.
The evaluation was conducted for three methods: ConceptGraphs [1], OpenScene [3], and a very recent BBQ [2]. We used 2D fused features for OpenScene, as this mode demonstrated the best performance. These methods employ different approaches, so we can perform benchmarking on a wider spectrum.
For more fair comparison between the methods, we replaced LLaVa-7b-v0 [22] and GPT4 [23] originally suggested by ConceptGraphs by LLaVa1.6-7b [24] and LLaMa3.1-8B [25], also used by BBQ. While LLM and LVLM play an important role in building a semantic graph, it is not a part of reported contributions, and most approaches use it off-the-shelf. Moreover, we wanted to verify performance given a more compact model.
Obtained results for segmentation quality metrics for all considered methods and scenes are presented in Tables III, while their relative change under different test conditions is illustrated on a ReplicaCAD dataset by Fig. 5. Final metrics are computed differently for each dataset: for ReplicaCAD, we calculate an overall confusion matrix since it provides a consistent closed vocabulary across all scenes, while for HM3D, we average metrics across individual scenes due to the large number of unique labels and potential label inconsistencies for identical objects across different scenes.
Semantic graph quality assessment has been performed according to the described methodology. Fig 6 illustrates VQA accuracy achieved by BBQ and ConceptGraphs methods for different categories of semantic queries. We did not evaluate OpenScene because it does not construct a scene graph.
For BBQ [2], questions related to object relations were not considered, as its scene graph construction is query-dependent. Instead, the evaluation focused on assessing the correctness of object presence and attributes, ensuring that the retrieved object set aligns with the given question.
In addition, we measure time consumption using NVIDIA GeForce RTX 4090 for 3D scene representation in frames per second (FPS), which is shown in Tables III.
@@ V-C Results Analysis
Collecting mAcc and f-mIoU metrics from testing ConceptGraphs, BBQ, and OpenScene on scenes in ReplicaCAD and HM3D reveals key trade-offs. f-mIoU is weighted by class frequency, so voluminous objects like walls and floors matter more, while mAcc gives equal weight to all objects.
On ReplicaCAD, OpenScene achieves the highest f-mIoU, performing best on voluminous objects. ConceptGraphs achieves the highest mAcc, indicating strength in recognizing more rare classes. BBQ performs steadily across both metrics but doesn’t lead in either. Under the velocity setting, OpenScene improves on both metrics, likely due to better filtering with fewer points. Under camera light, both BBQ and ConceptGraphs achieve higher scores in mAcc, likely because the direct lighting emphasizes objects, boosting detection. Also, ConceptGraphs is most sensitive to light conditions. On HM3D, performance drops across the board, but OpenScene shows the best performance.
Figure 6 presents the answering accuracy across different question categories and scene modifications. Binary existence-based questions exhibit high accuracy, indicating that models do not hallucinate non-existent objects. Overall, when binary questions achieve accuracy above 50–60%, it confirms non-random response patterns.
@@
The highest accuracy is observed under nominal conditions — the baseline for ReplicaCAD and nominal lighting for HM3D. Variations in lighting and velocity reduce accuracy, especially in object attributes and relationships questions. Attribute-based questions often fail due to missing material and property details in scene graph captions, leading to incomplete LLM responses.
Taking into account the analysis of metrics, it was also possible to detect corner cases (Fig. 7) under variable lighting conditions, which can lead to a decrease in accuracy. When using camera light, shadows of objects can be counted as part of the object. Dynamic lighting causes incorrect segmentation of uniform surfaces, such as floors, due to illumination differences. Also, when using only the camera light source, a noisy point cloud was obtained. The tested methods show sensitivity to objects with a similar description (for example, “wall cabinet” and “base cabinet”). Also, when assessing the quality of semantic segmentation, it turned out that the methods show worse results if we use a ‘_’ instead of a space in the class name.
There is also an object merging issue that affects the answers to Measurement-based questions. For example, in ReplicaCAD, both methods mark each step of the stairs as a separate staircase, e.g.:
@@ VI Conclusion
In this paper, we target a systematic evaluation of open semantic mapping algorithms, in particular their robustness in dynamic environments, e.g., under varying indoor lighting conditions. To accomplish the task, we introduced a dynamically configurable and highly automated LLM/LVLM-powered benchmarking methodology.
Our benchmarking procedure enabled a comprehensive analysis of leading methods, such as OpenScene [3], BBQ [2], and Concept Graph [1], highlighting the impact of lighting variations on semantic segmentation accuracy and constructed knowledge graphs quality.
@@
Our solution offers a flexible framework that allows easy extension to evaluate additional capabilities of OSM approaches. OSMa-Bench’s adaptability allows it to generate scenes with dynamic objects, evaluate visual grounding, and assess spatial reasoning by expanding the question categories within the existing pipeline. We hope that our code implementation will be useful for further research.